# GeoGraph: Geometric and Graph-based Ensemble Descriptors for Intrinsically Disordered Proteins

### v1

@@ 1 Introduction
Proteins are the cell’s molecular machines: sequence-encoded biopolymers which catalyze reactions, regulate processes, and shape cellular architecture. Recent years have witnessed a paradigm shift in protein modelling, driven by advances in experimental techniques and the maturation of deep learning. In particular, the rapid growth of high-throughput sequencing has been pivotal [34]. On the one hand it has enabled language-modelling approaches, especially Masked Language Modelling (MLM), to learn the statistical patterns of evolution directly from vast, unannotated sequence databases [28, 22]. On the other, Multiple Sequence Alignments (MSAs), coupled with decades of structure determination experiments [4], underpin deep learning models like AlphaFold [20] and RosettaFold [3], which now achieve near-experimental accuracy for a broad class of structured proteins.
@@
With static structures largely tractable, the frontier of computational structural biology is advancing toward a more fundamental problem: modelling the full conformational ensemble – the Boltzmann distribution of conformations under physiological solution conditions. To frame this challenge, we can identify three regimes along the structural order-disorder continuum: (i) proteins that adopt a single, highly stable fold; (ii) dynamic proteins that interconvert among a few metastable states; and (iii) Intrinsically Disordered Proteins (IDPs), which manifest a broad, heterogeneous set of rapidly fluctuating conformations [39, 35]. The first regime is where models trained on protein crystal structures excel. The second is well-captured by Markov State Models (MSM), which characterise the ensemble by the populations of metastable states and the kinetic rates between them, typically inferred from long Molecular Dynamics (MD) simulations [27, 8, 15]. The third regime of IDPs is, however, particularly challenging, and provides the focus for this work. Beyond the inherent complexity of modelling a heterogeneous ensemble, these proteins also face significant experimental and evolutionary hurdles. Experimentally, obtaining data is laborious, and their dynamic nature means measurements are typically averaged across the entire ensemble and/or over time. Evolutionarily, they exhibit poor sequence conservation, a characteristic thought to derive from the lack of a stable structure required to maintain function [7].
@@
A recent line of work aims to use deep generative models, especially diffusion models, to map sequence directly to a full conformational distribution [21, 18, 16, 41]. While useful, this strategy faces practical and statistical hurdles: generating, storing, and analyzing thousands of conformers per protein is expensive, and for many downstream tasks such high-dimensional stochastic detail can obscure the underlying biophysical signal. From a statistical-physics perspective, fluctuations faster than the timescale of interest are effectively marginalized as entropy, making the explicit modelling of fine-grained, high-frequency detail counterproductive. Indeed timescale separation underpins MSM coarse-graining, which emphasizes slow, kinetically relevant transitions between states rather than the noisy internal motions within them [8, 15].
@@
Here we take a different approach: rather than modelling entire ensembles explicitly, we model their aggregate properties directly. Specifically, we propose to extract essential biophysical content of an IDP ensemble from the statistics of its transient residue–residue contacts [6, 2]. The power of this approach has recently been demonstrated by WARIO [12], which uses contact-based descriptors to cluster simulation trajectories of IDPs into structurally coherent states. Our work leverages this same insight for a different purpose: instead of post-hoc analysis of a single ensemble, our aim is high-throughput prediction directly from sequence. To achieve this, we convert conformations from simulation into residue-level contact-map graphs, compute a diverse set of graph-theoretic descriptors, and use their ensemble-averaged values as the direct prediction targets for our model. This approach acts as a deliberate information bottleneck, filtering high-frequency fluctuations while preserving the stable signature of the ensemble.
@@ 2 Related work
Our work builds on several research threads at the intersection of machine learning and protein science. One major line of work uses deep generative models to sample full conformational ensembles from sequence, for both general proteins [21, 18, 9] and specifically for IDPs [26, 17, 16, 41]. While powerful, these methods can be computationally expensive, and general-purpose models often rely on co-evolutionary signals from MSAs that are absent in IDPs. An alternative approach, more aligned with our own, is to predict ensemble-averaged aggregate properties directly from sequence. For example, ALBATROSS [24] predicts five global geometric properties of IDPs, while IDP-BERT [25] fine-tunes a protein language model for similar tasks. We extend this strategy by introducing GeoGraph, a model that learns a richer representation by predicting a diverse set of residue-level geometric and graph-theoretic descriptors. This method is inspired by the long history of using residue contact networks to analyze protein structure and stability [6, 2] and is complementary to the recent method WARIO [12] which uses contact maps for post-hoc characterization of individual IDP ensembles.
@@ 3 GeoGraph
Generating the vast simulation data required for deep learning is computationally prohibitive with high-fidelity all-atom force fields. We therefore use CALVADOS-2 [33], a state-of-the-art one-bead-per-residue coarse-grained force field designed for IDPs, and experimentally validated on SAXS and FRET measurements. While this approach sacrifices fine-grained details, its design is based on an effective description of non-bonded interactions, which enables it to excel at capturing transient residue-residue contact patterns.
@@
We consider two flavours of descriptors, which we refer to as geometric and graph-based. The geometric descriptors are commonly-employed sequence-level measures of IDP conformational ensembles: end-to-end distance ($R_{e}$), radius of gyration ($R_{g}$), asphericity ($\Delta$), and the Flory scaling exponent ($\nu$) and prefactor ($A_{0}$). Both $R_{g}$ and $R_{e}$ can be experimentally determined, and in turn used to determine the Flory prefactor and exponent [1]. Small Angle X-ray Scattering (SAXS) yields the ensemble-averaged radius of gyration $\langle R_{g}\rangle$, whereas Fluorescence resonance energy transfer (FRET) spectroscopy yields $\langle R_{e}\rangle$, or even $R_{e}$ distributions in the case of single-molecule FRET [14].
@@
We train and evaluate our models on the Human–IDRome dataset [32], containing simulated conformational ensembles for 28,058 intrinsically disordered regions from the human proteome. This is the largest publicly available dataset of its kind, which makes it ideal for benchmarking differing approaches. The ensembles were generated using the CALVADOS-2 coarse-grained force field, with each sequence represented by 1,000 weakly correlated conformational frames sampled from the simulation trajectory [32]. We partition the dataset using a 80/10/10 split based on sequence similarity, see Appendix A.3.
@@ 4 Evaluation
We compare against two leading IDP methods: STARLING [26], a generative diffusion model, and ALBATROSS [24], an RNN-based direct predictor. Since the original models were trained on data generated with a different force field, we retrained them on our dataset for a fair comparison. See Appendix D for further details.
We also compare the sequence-to-sequence backbone against Protein Language Model (PLM) embeddings. We use ESM-2 and test the model in two settings. Firstly, we used the general-purpose pre-trained embeddings of the 8M and 150M models. Secondly, we curated a dataset of 30 million IDP sequences, which we refer to as IDP-Euka-90, and used this to fine-tune two corresponding versions of ESM-2, IDP-ESM2-8M and IDP-ESM2-150M, see Appendix B for further details. In both cases, we freeze the backbone model and train a sequence-level prediction head for predicting the geometric features – as we did for the Graph model above.
Finally, we attempted to evaluate BioEmu, a large-scale general-purpose ensemble emulator [21] which uses a diffusion model to generate conformational ensembles conditioned on the MSA of a sequence. Due to computational constraints, we were not able to generate sufficiently large ensembles with BioEmu on our test set to make a fair comparison. In a small experiment where we generated 1000 conformers/sequence for 100 randomly-sampled test sequences, we observed very poor performance ($R^{2}<0$ for all features), which is consistent with recent work evaluating BioEmu for IDPs [29], and may be explained by the poor sequence conservation of IDPs.
@@ 5 Discussion
As shown in Table 1, GeoGraph achieves highly competitive performance against leading methods for IDP ensemble property prediction, in particular on the more complex shape descriptors $(\Delta,\nu,A_{0})$. Critically, our model predicts these descriptors several orders of magnitude faster than it takes to run the CALVADOS-2 simulator that it emulates: GeoGraph can process the entire test set of 2,388 sequences in approximately 1 second on a single GPU (H100), whereas simulation of these ensembles on Google Colab takes on the order of 10 days [32].
The source of GeoGraph’s strong performance can be seen by comparing our model variants: the Geo model, which trains on geometric features alone, and the Graph model, which learns representations solely from graph topology, and is evaluated on the GeoHead trained on these. Both variants perform on par with each other, demonstrating that the rich biophysical information in the contact-map topology is sufficient to create representations as powerful as those learned by direct optimization. The value of these learned representations is confirmed by the far superior performance of Graph relative to our Geo-zero baseline, which lacks this contextual learning. Crucially, the main GeoGraph model outperforms both specialized variants, demonstrating a clear synergistic effect. This supports our central hypothesis: the auxiliary task of predicting contact map characteristics is a highly beneficial component for extracting transferable representations from MD simulation data. Furthermore, the ablation in Table 2 reveals that the context-aware, residue-level graph features are the primary drivers of this learning.
@@
For the comparison against PLM embeddings, the most direct reference point is with our Graph $\to$ GeoHead model. We see that our simulation-informed embeddings provide a significantly stronger predictive signal for geometric properties, even after fine-tuning ESM-2 on a large corpus of IDP sequences. While a superior performance may be expected when the training objective aligns with the evaluation task, the magnitude of the difference underscores a key limitation of protein language models when applied to IDPs: their reliance on evolutionary patterns, which serve as a noisy and incomplete proxy for the physical properties that govern dynamic ensembles [7].
@@ 6 Conclusion
In this work we introduce GeoGraph, a sequence-to-sequence model trained to predict aggregate properties of IDP conformational ensembles. It achieves this by first featurizing individual conformations from MD simulations into contact-graph topologies, and then learning to predict the ensemble average of these features, at both a residue- and sequence-level. Our evaluation demonstrates that this approach not only achieves highly competitive performance on benchmark tasks but also yields embeddings that are more effective for predicting key experimentally relevant properties than existing methods. Our trained GeoGraph and IDP-ESM2 models, along with the IDP-Euka-90 training dataset, will be publicly released.
@@ A.1 GeoGraph
GeoGraph is a sequence-to-sequence model that maps a protein’s amino-acid sequence to feature vectors describing aggregate physical properties at both the sequence- and residue-level. The backbone is a transformer encoder [36], chosen for its ability to capture long-range dependencies and produce context-rich embeddings. We build on the Hugging Face implementation of ESM-2 [22, 38], which uses Pre-Layer Normalization (Pre-LN) [40] and Rotary Position Embeddings (RoPE) [31].
@@ A.2 Features
We consider two flavours of descriptors, which we refer to as geometric and graph-based. The geometric descriptors all sequence-level features, while for the graph-based descriptors we consider both sequence- and residue-level features. Graph features are computed using python’s NetworkX package (default settings) [13], and for training our models we standardise all target features to have zero mean and unit variance.
@@
Graph, residue-level: Here we included degree centrality (local contact density), betweenness centrality (bridging propensity), harmonic centrality (inverse-distance reachability), PageRank [5], core number, local clustering coefficient, and as well as an in-largest-connected-component indicator.
@@ A.3 Human–IDRome dataset
We partitioned the Human–IDRome dataset [32] based on sequence similarity into 80/10/10 splits for training, validation, and testing. To ensure fair comparison with prior work, this split was performed using MMseqs2 [30] with parameters (min_seq_id=0.7, coverage=0.8, cov_mode=1), identical to the parameters used by STARLING [26]. Additionally, we filtered the dataset to sequences with a maximum length of 256 residues.
@@ Appendix B IDP-ESM
For training our fine-tuned versions of ESM-2, IDP-ESM2-8M and IDP-ESM2-150M, we curated a large dataset of biological IDP sequences, which we call IDP-Euka-90. As suggested in the Metapredict V3 paper [23], eukaryotes have significantly more disordered regions than bacteria and euryarchaeota: we hence decided to focus on eukaryotes to extract IDRs. We downloaded all 2764 eukaryota proteomes from UniProt and ran Metapredict V3 command metapredict-predict-idrs [11] with default disorder threshold of 0.5 on each one of them. We removed sequences shorter than 30 amino acids and clustered the dataset with mmseqs2 linclust command, with minimum sequence identity threshold of 0.9, 0.8 coverage in coverage mode 1. This pipeline produced an IDP dataset consisting of 30,337,340 sequences.

### v2

@@ 1 Introduction
Proteins are the cell’s molecular machines: sequence-encoded biopolymers which catalyze reactions, regulate processes, and shape cellular architecture. Recent years have witnessed a paradigm shift in protein modelling, driven by advances in experimental techniques and the maturation of deep learning. In particular, the rapid growth of high-throughput sequencing has been pivotal [theuniprotconsortium2025]. On the one hand it has enabled language-modelling approaches, especially Masked Language Modelling (MLM), to learn the statistical patterns of evolution directly from vast, unannotated sequence databases [Rives2021, Lin2023ESMFold]. On the other, Multiple Sequence Alignments (MSAs), coupled with decades of structure determination experiments [Berman2000], underpin deep learning models like AlphaFold [Jumper2021] and RosettaFold [baek2021rosettafold], which now achieve near-experimental accuracy for a broad class of structured proteins.
@@
With static structures largely tractable, the frontier of computational structural biology is advancing toward a more fundamental problem: modelling the full conformational ensemble – the Boltzmann distribution of conformations under physiological solution conditions. To frame this challenge, we can identify three regimes along the structural order-disorder continuum: (i) proteins that adopt a single, highly stable fold; (ii) dynamic proteins that interconvert among a few metastable states; and (iii) Intrinsically Disordered Proteins (IDPs), which manifest a broad, heterogeneous set of rapidly fluctuating conformations [DysonWright2005, VanderLee2014]. The first regime is where models trained on protein crystal structures excel. The second is well-captured by Markov State Models (MSM), which characterise the ensemble by the populations of metastable states and the kinetic rates between them, typically inferred from long Molecular Dynamics (MD) simulations [Prinz2011, Chodera2014, Husic2018]. The third regime of IDPs is, however, particularly challenging, and provides the focus for this work. Beyond the inherent complexity of modelling a heterogeneous ensemble, these proteins also face significant experimental and evolutionary hurdles. Experimentally, obtaining data is laborious, and their dynamic nature means measurements are typically averaged across the entire ensemble and/or over time. Evolutionarily, they exhibit poor sequence conservation, a characteristic thought to derive from the lack of a stable structure required to maintain function [brown2011idpevolution].
@@
A recent line of work aims to use deep generative models, especially diffusion models, to map sequence directly to a full conformational distribution [lewis2025bioemu, jin2025p2dflow, janson2024idpsam, zhu2024idpfold]. While useful, this strategy faces practical and statistical hurdles: generating, storing, and analyzing thousands of conformers per protein is expensive, and for many downstream tasks such high-dimensional stochastic detail can obscure the underlying biophysical signal. From a statistical physics perspective, fluctuations faster than the timescale of interest are effectively marginalized as entropy, making the explicit modelling of fine-grained, high-frequency detail counterproductive. Indeed timescale separation underpins MSM coarse-graining, which emphasizes slow, kinetically relevant transitions between metastable states, rather than the noisy internal motions within them [Chodera2014, Husic2018].
@@
Here we take a different approach: rather than modelling entire ensembles explicitly, we model their aggregate properties directly. Specifically, we propose to extract essential biophysical content of an IDP ensemble from the statistics of its transient residue–residue contacts [brinda2005, amitai2004]. The power of this approach has recently been demonstrated by WARIO [gonzalezdelgado2024wario], which uses contact-based descriptors to cluster simulation trajectories of IDPs into structurally coherent states. Our work leverages this same insight for a different purpose: instead of post-hoc analysis of a single ensemble, our aim is high-throughput prediction directly from sequence. To achieve this, we convert conformations from simulation into residue-level contact-map graphs, compute a diverse set of graph-theoretic descriptors, and use their ensemble-averaged values as the direct prediction targets for our model. This approach acts as a deliberate information bottleneck, filtering high-frequency fluctuations while preserving the stable signature of the ensemble.
@@
We make our code, pre-trained models and datasets publicly available at https://github.com/instadeepai/GeoGraph.
@@ 2 Related work
Our work builds on several research threads at the intersection of machine learning and protein science. One major line of work uses deep generative models to sample full conformational ensembles from sequence, for both general proteins [lewis2025bioemu, jin2025p2dflow, daigavane2025jamun] and specifically for IDPs [novak2025, janson2023idpgan, janson2024idpsam, zhu2024idpfold]. While powerful, these methods can be computationally expensive, and general-purpose models often rely on co-evolutionary signals from MSAs that are absent in IDPs. An alternative approach, more aligned with our own, is to predict ensemble-averaged aggregate properties directly from sequence. For example, ALBATROSS [lotthammer2024] predicts five global geometric properties of IDPs, while IDP-BERT [mollaei2024idpbert] fine-tunes a protein language model for similar tasks. We extend this strategy by introducing GeoGraph, a model that learns a richer representation by predicting a diverse set of residue-level geometric and graph-theoretic descriptors. This method is inspired by the long history of using residue contact networks to analyze protein structure and stability [brinda2005, amitai2004] and is complementary to the recent method WARIO [gonzalezdelgado2024wario] which uses contact maps for post-hoc characterization of individual IDP ensembles.
@@ 3 GeoGraph
Generating the vast simulation data required for deep learning is computationally prohibitive with high-fidelity all-atom force fields. We therefore use CALVADOS-2 [Tesei2023CALVADOS2], a state-of-the-art one-bead-per-residue coarse-grained force field designed for IDPs, and experimentally validated on SAXS and FRET measurements. While this approach sacrifices fine-grained detail, its design is based on an effective description of non-bonded interactions, which enables it to excel at capturing transient residue-residue contact patterns.
@@
We consider two flavours of descriptors, which we refer to as geometric and graph-based. The geometric descriptors are commonly-employed sequence-level measures of IDP conformational ensembles: end-to-end distance ($R_{e}$), radius of gyration ($R_{g}$), asphericity ($\Delta$), and the Flory scaling exponent ($\nu$) and prefactor ($A_{0}$). Both $R_{g}$ and $R_{e}$ can be experimentally determined, and in turn used to determine the Flory prefactor and exponent [ahmed2019]. Small Angle X-ray Scattering (SAXS) yields the ensemble-averaged radius of gyration $\langle R_{g}\rangle$, whereas Fluorescence resonance energy transfer (FRET) spectroscopy yields $\langle R_{e}\rangle$, or even $R_{e}$ distributions in the case of single-molecule FRET [hofmann2012].
@@
We train and evaluate our models on the Human–IDRome dataset [tesei2024idrome], containing simulated conformational ensembles for 28,058 intrinsically disordered regions from the human proteome. This is the largest publicly available dataset of its kind, which makes it ideal for benchmarking differing approaches. The ensembles were generated using the CALVADOS-2 coarse-grained force field, with each sequence represented by 1,000 weakly correlated conformational frames sampled from the simulation trajectory [tesei2024idrome]. We partition the dataset using a 80/10/10 split based on sequence similarity, see Appendix A.3.
@@ 4 Evaluation
We compare against two leading IDP methods: STARLING [novak2025], a generative diffusion model, and ALBATROSS [lotthammer2024], an RNN-based direct predictor. Since the original models were trained on data generated with a different force field, we retrained them on our dataset for a fair comparison. See Appendix D for further details.
In their analysis of the Human–IDRome dataset [tesei2024idrome], the authors trained a support vector regression model on a manually chosen set of sequence features for modelling the Flory scaling exponent $\nu$. For comparison we evaluated this on our test set, and find it achieves an $R^{2}$ score of 0.765. The author’s also quantified the simulation (convergence) error for $\nu$ as $\sim 0.01$. For GeoGraph we find the corresponding RSME as 0.011, which provides a strong sanity check that the model is learning the physical signal right up to the limit of the data’s precision.
We next compare the sequence-to-sequence backbone against Protein Language Model (PLM) embeddings. We use ESM-2 [Lin2023ESMFold] and test the model in two settings. Firstly, we used the general-purpose pre-trained embeddings of the 8M and 150M models. Secondly, we curated a dataset of 30 million IDP sequences, which we refer to as IDP-Euka-90, and used this to fine-tune two corresponding versions of ESM-2, IDP-ESM2-8M and IDP-ESM2-150M, see Appendix B for further details. In both cases, we freeze the backbone model and train a sequence-level prediction head for predicting the geometric features – as we did for the Graph model above.
Finally, we attempted to evaluate BioEmu, a large-scale general-purpose ensemble emulator [lewis2025bioemu] which uses a diffusion model to generate conformational ensembles conditioned on the MSA of a sequence. Due to computational constraints, we were not able to generate sufficiently large ensembles with BioEmu on our test set to make a fair comparison. In a small experiment where we generated 1000 conformers/sequence for 100 randomly-sampled test sequences, we observed very poor performance ($R^{2}<0$ for all features), which is consistent with recent work evaluating BioEmu for IDPs [schnapka2025atomic], and may be explained by the poor sequence conservation of IDPs.
@@ 5 Discussion
As shown in Table 1, GeoGraph achieves highly competitive performance against leading methods for IDP ensemble property prediction, in particular on the more complex shape descriptors $(\Delta,\nu,A_{0})$. Critically, our model predicts these descriptors several orders of magnitude faster than it takes to run the CALVADOS-2 simulator that it emulates: GeoGraph can process the entire test set of 2,388 sequences in approximately 1 second on a single GPU (H100), whereas simulation of these ensembles on Google Colab takes on the order of 10 days [tesei2024idrome].
The source of GeoGraph’s strong performance can be seen by comparing our model variants: the Geo model, which trains on geometric features alone, and the Graph model, which learns representations solely from graph topology (and is evaluated on the GeoHead trained on these). Both variants perform on par with each other, demonstrating that the rich biophysical information in the contact-map topology is sufficient to create representations as powerful as those learned by direct optimization. The value of these learned representations is confirmed by the far superior performance of Graph relative to our Geo-zero baseline, which lacks this contextual learning. Crucially, the main GeoGraph model outperforms both specialized variants, demonstrating a clear synergistic effect. This supports our central hypothesis: the auxiliary task of predicting contact map characteristics is a highly beneficial component for extracting transferable representations from MD simulation data. Furthermore, the ablation in Table 2 reveals that the context-aware, residue-level graph features are the primary drivers of this learning.
@@
For the comparison against PLM embeddings, the most direct reference point is with our Graph $\to$ GeoHead model. We see that our simulation-informed embeddings provide a significantly stronger predictive signal for geometric properties, even after fine-tuning ESM-2 on a large corpus of IDP sequences. While a superior performance may be expected when the training objective aligns with the evaluation task, the magnitude of the difference underscores a key limitation of protein language models when applied to IDPs: their reliance on evolutionary patterns, which serve as a noisy and incomplete proxy for the physical properties that govern dynamic ensembles [brown2011idpevolution].
@@ 6 Conclusion
In this work we introduce GeoGraph, a sequence-to-sequence model trained to predict aggregate properties of IDP conformational ensembles. It achieves this by first featurizing individual conformations from MD simulations into contact-graph topologies, and then learning to predict the ensemble average of these features, at both a residue- and sequence-level. Our evaluation demonstrates that this approach not only achieves highly competitive performance on benchmark tasks but also yields embeddings that are more effective for predicting key experimentally relevant properties than existing methods. Our trained GeoGraph and IDP-ESM2 models, along with the IDP-Euka-90 training dataset are available at https://github.com/instadeepai/GeoGraph.
@@ A.1 GeoGraph
GeoGraph is a sequence-to-sequence model that maps a protein’s amino-acid sequence to feature vectors describing aggregate physical properties at both the sequence- and residue-level. The backbone is a transformer encoder [vaswani2017attention], chosen for its ability to capture long-range dependencies and produce context-rich embeddings. We build on the Hugging Face implementation of ESM-2 [Lin2023ESMFold, Wolf2020Transformers], which uses Pre-Layer Normalization (Pre-LN) [Xiong2020PreLN] and Rotary Position Embeddings (RoPE) [Su2021RoPE].
@@ A.2 Features
We consider two flavours of descriptors, which we refer to as geometric and graph-based. The geometric descriptors all sequence-level features, while for the graph-based descriptors we consider both sequence- and residue-level features. Graph features are computed using python’s NetworkX package (default settings) [hagberg2008], and for training our models we standardise all target features to have zero mean and unit variance.
@@
Graph, residue-level: Here we included degree centrality (local contact density), betweenness centrality (bridging propensity), harmonic centrality (inverse-distance reachability), PageRank [brin1998pagerank], core number, local clustering coefficient, and as well as an in-largest-connected-component indicator.
@@ A.3 Human–IDRome dataset
We partitioned the Human–IDRome dataset [tesei2024idrome] based on sequence similarity into 80/10/10 splits for training, validation, and testing. To ensure fair comparison with prior work, this split was performed using MMseqs2 [steinegger2017mmseqs2] with parameters (min_seq_id=0.7, coverage=0.8, cov_mode=1), identical to the parameters used by STARLING [novak2025]. Additionally, we filtered the dataset to sequences with a maximum length of 256 residues.
@@ Appendix B IDP-ESM
For training our fine-tuned versions of ESM-2, IDP-ESM2-8M and IDP-ESM2-150M, we curated a large dataset of biological IDP sequences, which we call IDP-Euka-90. As suggested in the Metapredict V3 paper [Lotthammer2024.11.05.622168], eukaryotes have significantly more disordered regions than bacteria and euryarchaeota: we hence decided to focus on eukaryotes to extract IDRs. We downloaded all 2764 eukaryota proteomes from UniProt and ran Metapredict V3 command metapredict-predict-idrs [emenecker2021metapredict] with default disorder threshold of 0.5 on each one of them. We removed sequences shorter than 30 amino acids and clustered the dataset with mmseqs2 linclust command, with minimum sequence identity threshold of 0.9, 0.8 coverage in coverage mode 1. This pipeline produced an IDP dataset consisting of 30,337,340 sequences.